OpenStudy (anonymous):

Solve 6 + 5z = -12(3 – z). Please show all of your work.

OpenStudy (anonymous):

\[6 + 5z = -12(3 – z)\]\[6+5z=-36+12z\]\[6+36=12z-5z\]\[42=7z\]\[z=6\]